# Personalizing Student-Agent Interactions Using Log-Contextualized Retrieval Augmented Generation (RAG)

### v1

@@ 4.2 Copa (RQ2)
RQ2 asked, “How do students perceive the epistemic value of Copa’s guidance, and to what extent does Copa support their critical thinking?” We addressed this question through qualitative analysis of agent interactions, Miro boards, class discussions, and survey responses from the classroom and focus group studies (see Section 3.2). Building on findings from RQ1, we refined LC-RAG in two ways. First, we used a condensed knowledge base consisting of 15 task-specific chunks (e.g., kinematic equations, computing concepts like initializing and updating variables). Second, we adopted an agentic approach (i.e., allowing the LLM to analyze data, reason, and act) to integrating log and text data. Rather than relying on predefined task context categories, we provided the LLM with the current state of the students’ computational model. During summarization, the LLM identified the problem students expressed, diagnosed its cause by comparing the student model to an expert version, and recommended relevant domain knowledge to retrieve. This recommendation was then embedded for semantic search. Copa’s summarization and interface prompts were iteratively refined by this paper’s authors and provided in the Supplementary Materials.
Overall, students found Copa to be epistemically valuable. All six focus group participants reported that their interactions with Copa supported their problem-solving and that the information it provided was accurate. This sentiment was echoed in classroom discussions and on the Miro boards, where many students emphasized how Copa helped them better understand the course material (“It felt like Copa was in the room with us the entire time trying to teach us”;“Has a step-by-step process that is easy to follow.”;“Provided a reflection after we asked the questions”). Several students also highlighted Copa’s positive tone as a source of motivation and encouragement (“…polite and encouraging feedback”,“… praising us when we got a correct answer”,“My favorite part about COPA was its motivative language”). One student, who had struggled with the C2STEM Truck Task two years earlier when no agent was available, shared that she felt “a lot more confident” in her abilities and “got a lot further on my own this time around”.
Copa supported students’ critical thinking and epistemic decision-making (“…helped me critically think…”), enabling them to grasp domain concepts they had previously struggled to understand. In one instance, one group admitted to Copa that they “…have no idea…” how to calculate the lookahead distance — i.e., the distance required for the truck to decelerate and stop at a stop sign. Recognizing their need for guidance, Copa retrieved and presented a relevant kinematic equation from the knowledge base using LC-RAG. With this information, the students realized they needed to calculate the truck’s displacement, gaining a clearer understanding of its role in both the equation and the deceleration phase of the Truck Task.
@@ 5 Discussion and Conclusions
Methodologically, our RQ2 findings suggest that agentic reasoning enhances personalized student support. Through characterizing student difficulties and recommending domain knowledge — based on students’ verbalized needs and grounded in their actions within the C2STEM environment — Copa generates embeddings that align closely with the knowledge base, enabling effective RAG retrieval and more contextualized interactions. This process introduces a transparent reasoning chain: model outputs are grounded in retrieved content, which are tied to the LLM’s reasoning and can be traced back to student actions and expressed problems. This structure not only supports effective personalization but also enhances explainability, offering insight into both what the agent recommends and why. We hypothesize that additional conversational agent components (e.g., dialogue managers), could similarly benefit from agentic reasoning, positioning agentic AIED as a compelling direction for future research.
@@
More broadly, this underscores the importance of grounding LLM-powered agents in established learning theories. The accessibility and ease of use of LLMs have led to a surge in pedagogical agents often deployed without a clear theoretical foundation — marking a departure from pre-LLM intelligent tutoring systems, which were typically tightly aligned with cognitive and learning science principles. Recent AIED research has raised concerns about the risks of designing systems that overlook the pedagogical foundations of agent feedback [37]. As LLMs become increasingly central to AIED, theory-driven design is essential to ensure these systems do more than mimic helpfulness — they must reflect how students actually learn. Grounding intelligent systems in learning theory is also critical for building trust among practitioners, who require confidence that AI interventions — particularly GenAI ones — are pedagogically sound, not just technologically advanced.

### v2

@@ 4.2 Copa (RQ2)
We answer RQ2 by qualitatively analyzing agent interactions, Miro boards, class discussions, and survey responses from the classroom and focus group studies (see Section 3.3.2). Building on findings from RQ1, we refined LC-RAG in two ways. First, we used a condensed knowledge base consisting of 15 task-specific chunks (e.g., kinematic equations, computing concepts like initializing and updating variables). Second, we adopted an agentic approach (i.e., allowing the LLM to analyze data, reason, and act) to integrating log and text data. Rather than relying on predefined task context categories, we provided the LLM with the current state of the students’ computational model. During summarization, the LLM identified the problem students expressed, diagnosed its cause by comparing the student model to an expert version, and recommended relevant domain knowledge to retrieve. This recommendation was then embedded for semantic search. Copa’s summarization and interface prompts were iteratively refined by this paper’s authors and are provided in the Supplementary Materials.
Overall, students found Copa to be epistemically valuable. All focus group participants reported that their interactions with Copa supported their problem-solving and the agent’s information was accurate. This sentiment was echoed in classroom discussions and on the Miro boards, where many students emphasized how Copa helped them better understand the course material (“It felt like Copa was in the room with us the entire time trying to teach us”;“Has a step-by-step process that is easy to follow.”;“Provided a reflection after we asked the questions”). Several students also highlighted Copa’s positive tone as a source of motivation and encouragement (“…polite and encouraging feedback”,“… praising us when we got a correct answer”,“My favorite part about COPA was its motivative language”). One student, who had struggled with the C2STEM Truck Task two years earlier when no agent was available, shared that she felt “a lot more confident” in her abilities and “got a lot further on [her] own this time around.”
Copa supported students’ critical thinking and epistemic decision-making (“…helped me critically think…”), enabling them to grasp domain concepts they had previously struggled to understand. In one instance, one group admitted to Copa that they “…have no idea…” how to calculate the lookahead distance — i.e., the distance required for the truck to decelerate and stop at the stop sign. Recognizing their need for guidance, Copa retrieved and presented a relevant kinematic equation from the knowledge base using LC-RAG. With this information, the students realized they needed to calculate the truck’s displacement, gaining a clearer understanding of its role in both the equation and the deceleration phase of the Truck Task.
@@ 5 Discussion and Conclusions
Methodologically, our RQ2 findings suggest that agentic reasoning enhances personalized student support. Through characterizing student difficulties and recommending domain knowledge — based on students’ verbalized needs and grounded in their actions within the C2STEM environment — Copa generates embeddings that align closely with the knowledge base, enabling effective RAG retrieval and more contextualized interactions. This process introduces a transparent reasoning chain: model outputs are grounded in retrieved content, which is tied to the LLM’s reasoning and can be traced back to student actions and expressed problems. This structure not only supports effective personalization but also enhances explainability, offering insight into both what the agent recommends and why. We hypothesize that (1) additional conversational agent components (e.g., dialogue managers), could similarly benefit from agentic reasoning; and (2) agentic reasoning will allow LLM-based agents to support students’ learning processes by deepening their conceptual understanding and promoting epistemic growth over time, positioning agentic AIED as a compelling direction for future research.
@@
More broadly, this underscores the importance of grounding LLM-powered agents in established learning theories. The accessibility and ease of use of LLMs have led to a surge in pedagogical agents often deployed without a clear theoretical foundation — marking a departure from pre-LLM intelligent tutoring systems, which were tightly aligned with cognitive and learning science principles. Recent AIED research has raised concerns about the risks of designing LLM systems that overlook the pedagogical foundations of agent feedback [25]. As LLMs become increasingly central to AIED, theory-driven design is essential to ensure these systems do more than mimic helpfulness — they must reflect how students actually learn. Grounding intelligent systems in learning theory is also critical for building trust among practitioners, who require confidence that AI interventions — particularly GenAI ones — are pedagogically sound, not just technologically advanced.